Circular No. 9035

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2009 F6 (YI-SWAN) Further to IAUC 9034, H. Yamaoka, Kyushu University, informs the

Central Bureau that he received an e-mailed report on March 28 of the discovery by Dae-am Yi (Yeongwol-kun, Gangwon-do, Korea) of a possible comet with a noticeable greenish coma of diameter ∼ 1′ (discovery observation tabulated below) on two 60-s CCD exposures taken about 80 seconds apart with a Canon 5D camera and a 90-mm f/2.8 camera lens. Orbital calculations suggest that this object is identical with C/2009 F6.

2009 UT  α_{2000}  δ_{2000}  Mag. Observer
Mar. 26.835 22^{h}39^{m}40^{s} +49^{o}49′08′′  12.5  Yi

The available precise astrometry, the following preliminary parabolic orbital elements from the precise astrometry only, and an ephemeris appear on MPEC 2009-G21.

T = 2009 May 8.056 TT  ω = 130.^{o}432
Ω = 278.927 }2000.0
q = 1.26991 AU  i = 85.724

Further visual total-magnitude and coma-diameter estimates: Apr. 6.16 UT, 8.1, 7′ (J. J. Gonzalez, Burgos, Spain, 25×100 binoculars); 6.91, 9.7, 2′ (V. Nevski, Vitebsk, Belarus, 0.30-m reflector); 7.06, 8.5, 4′ (B. H. Granslo, Fjellhamar, Norway, 0.20-m reflector; moonlight); 7.09, 8.5, 4′ (M. Reszelski, Szamotuly-Galowo, Poland, 15×70 binoculars).

S. Korotkiy and T. Kryachko, Moscow, report that their unfiltered CCD observations with a 30-cm telescope at Karachay-Cherkessia, Russia, on Apr. 7.06 show a bright asymmetrical coma of size 332′′ × 354′′ (major axis in p.a. 161^{o}) with an asymmetrical condensation of size 12′′.8 × 19′′.3 (major axis in p.a. 320^{o}).

2009 DD_{47}

Numerous attempts to find cometary appearance in the retrogradeorbiting object 2009 DD_{47} (cf. MPEC 2009-E12; T = 2008 Oct. 29.6206 TT, e = 0.976539, q = 2.242187 AU, ω = 315^{o}.1224, Ω = 173^{o}.8329, i = 107^{o}.8501, equinox 2000.0) have produced negative results. J. Young, Table Mountain Observatory, writes that observations taken on numerous nights in March by M. Hicks, H. Rhoades, and himself show only a starlike appearance.

2009 April 7  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2009 CBAT  Daniel W. E. Green